# Efficient Optimization of Variational Autoregressive Networks with Natural Gradient

### v1

@@ Abstract
Estimating free energy is a fundamental problem in statistical mechanics. Recently, machine-learning-based methods, particularly the variational autoregressive networks (VANs), have been proposed to minimize variational free energy and to approximate the Boltzmann distribution. VAN enjoys notable advantages, including exact computation of the normalized joint distribution and fast unbiased sampling, which are critical features often missing in Markov chain Monte Carlo algorithms. However, VAN also faces significant computational challenges. These include difficulties in the optimization of variational free energy in a complicated parameter space and slow convergence of learning. In this work, we introduce an optimization technique based on natural gradients to the VAN framework, namely ng-VAN, to enhance the learning efficiency and accuracy of the conventional VAN. The method has computational complexity cubic in the batch size rather than in the number of model parameters, hence it can be efficiently implemented for a large VAN model. We carried out extensive numerical experiments on the Sherrington-Kirkpatrick model and spin glasses on random graphs and illustrated that compared with the conventional VAN, ng-VAN significantly improves the accuracy in estimating free energy and converges much faster with shorter learning time. This allows extending the VAN framework’s applicability to challenging statistical mechanics problems that were previously not accessible.
@@ I Introduction
Recently, a prominent application of machine learning methods, variational autoregressive networks (VANs), have been introduced to solve statistical mechanics problems. In VAN, an autoregressive neural network is employed as a variational ansatz $q_{\theta}(\mathbf{s})$ to approximate the Boltzmann distribution ${p_{\mathrm{eq}}(\mathbf{s})}$ [1, 2, 3]. The learning process of VAN is no longer based on maximum likelihood but is akin to reinforcement learning, with the objective of minimizing the Kullback-Leibler divergence $D_{\mathrm{KL}}\left(q_{\theta}\middle\|p_{\mathrm{eq}}\right)$, which necessitates sampling from the variational distribution itself. The autoregressive characteristics of the model enjoy tractable likelihood and unbiased sampling, a significant advantage over traditional Markov chain Monte Carlo methods. Likewise, autoregressive models have also been employed to represent variational wave functions in studies of variational Monte Carlo (VMC) [4, 5, 6].
It is worth noting that the autoregressive neural networks belong to generative models in machine learning [7]. In recent years, generative models have emerged as a powerful framework for modeling high-dimensional complex data distributions [7, 8]. They utilize the chain rule of probabilities to parameterize the joint distribution of sequential data, such as text and images. Modern large language models, particularly those built on transformer architecture [9], exemplify autoregressive modeling by generating the conditional probability of the subsequent word given a preceding sequence of inputs. When trained on sufficiently large and diverse corpora, large language models can effectively approximate the underlying joint distribution of natural languages [10].
@@
Despite its advantages, VAN also faces considerable challenges, particularly regarding optimization difficulties associated with non-convex parameter landscapes. VAN typically relies on first-order gradient-based optimization algorithms. Although state-of-the-art optimizers such as Adam [11] incorporate momentum information and show superior performances compared to traditional stochastic gradient descent [12], they may still encounter difficulties in high-dimensional parameter space and exhibit slow convergence speed. This limitation necessitates the need for more advanced optimization methods. One promising approach for solving this issue is the natural gradient descent [13, 14, 15], which has already proven effective in reinforcement learning [16, 17] as well as in VMC scenarios [18, 19]. The fundamental concept behind the natural gradient is to account for the curvature of the loss function by employing a preconditioner to rescale the gradient updates. Geometrically, the natural gradient represents the steepest descent direction in the model’s distribution space rather than in the Euclidean parameter space. However, this method requires the inversion of the Fisher information matrix, a computational complexity of $\mathcal{O}({N_{p}^{3}})$ with $N_{p}$ variational parameters. In recent advancements in VMC [20, 21] and machine learning [22], researchers have proposed efficient algorithms for implementing natural gradient, especially in situations where the number of samples or batch size $N_{b}$ is significantly smaller than $N_{p}$. Rather than inverting the original $N_{p}\times N_{p}$ matrix, it only requires the inversion of a significantly smaller $N_{b}\times N_{b}$ matrix, referred to as neural tangent kernel, which shares the same spectrum properties as the Fisher information matrix. In this way, the natural gradient method can be applied to neural network models with up to $10^{6}$ parameters, which is far beyond traditional approaches.
@@
In this work, we propose the integration of the natural gradient into the VAN framework, namely ng-VAN, to enhance both learning efficiency and accuracy compared to the conventional VAN. In particular, we employ an efficient method for computing the inverse of the Fisher information matrix with computational complexity growing cubic in the batch size, rather than the number of parameters. Our results demonstrate that the natural gradient method can significantly accelerate the convergence of VAN during training and achieve more accurate estimated variational free energy compared to first-order optimizers such as Adam. The rest of the paper is organized as follows: in Section II, we first introduce VAN and its learning, followed by an introduction to the natural gradient method and how to implement it efficiently in the context of VAN learning. In Section III, we present the experimental results, and finally, in Section IV, we summarize our conclusions and discuss the potential applications of our methods.
@@ II.1 Variational autoregressive networks
The variational autoregressive networks (VANs) proposed in Ref. [1] employ a variational ansatz for the target Boltzmann distribution, parameterized by a set of learnable parameters $\theta=\{\theta_{i}\}_{i=1}^{N_{p}}$, where $N_{p}$ denotes the total number of parameters. This variational distribution, denoted as $q_{\theta}(\mathbf{s})$, obeys the autoregressive property, i.e., can be decomposed into the product of conditional probabilities:
@@
$F_{q}(\theta)=\expectationvalue{E(\mathbf{s})+\frac{1}{\beta}\ln q_{\theta}(\mathbf{s})}_{q},$ (3)
where $\expectationvalue{\cdots}_{q}$ denotes the expectation over the variational distribution $q_{\theta}$. Moreover, the variational free energy $F_{q}(\theta)$ serves as an upper bound for the exact free energy. A crucial component required for gradient-based optimization is the gradient of the variational free energy with respect to $\theta$. Denoting $R(\mathbf{s})=E(\mathbf{s})+(1/\beta)\ln q_{\theta}(\mathbf{s})$, the gradient can be expressed as
$\eta=\nabla_{\theta}F_{q}(\theta)=\expectationvalue{R(\mathbf{s})\nabla_{\theta}\ln q_{\theta}(\mathbf{s})}_{q}.$ (4)
A baseline-corrected form can also be employed for variance reduction, given by $\eta=\expectationvalue{(R(\mathbf{s})-b)\nabla_{\theta}\ln q_{\theta}(\mathbf{s})}_{q}$, where $b=\expectationvalue{R(\mathbf{s})}_{q}$ is a constant termed as baseline. This gradient expression is known as the score function gradient estimator [23], also referred to as the REINFORCE algorithm [24], a specific instance of the policy gradient method from reinforcement learning.
VAN and its related techniques have demonstrated state-of-the-art variational performance in solving statistical mechanics problems, including sampling and evaluation of physical observables for Ising and spin glass systems [1, 2, 3, 25, 26, 27, 28]. However, as illustrated in Section I, it also encounters several significant challenges. In contrast to the case where a dataset is available for maximum likelihood estimation, in the training of VAN, gradient estimation in each iteration requires drawing samples from the current variational distribution. Due to the sequential nature of autoregressive sampling, this process is highly time-consuming. Furthermore, gradients derived from samples based on a sub-optimal variational distribution from the preceding step may not provide sufficient information for the subsequent optimization, leading to poor overall performance.
@@ II.2 Natural gradient descent
Here we present a concise and self-contained derivation of the natural gradient within the framework of optimizing VAN. The natural gradient can be derived by identifying the update $\delta\theta$ that minimizes the variational free energy $F_{q}(\theta+\delta\theta)$, subject to the constraint that the KL divergence between the current variational distribution $q_{\theta}$ and the updated distribution $q_{\theta+\delta\theta}$ is equal to a specified value $\epsilon$. Mathematically, this is formulated as:
@@
The Lagrange function for this optimization problem is given by:
@@
The KL divergence can be approximated by $D_{\mathrm{KL}}\left(q_{\theta}\middle\|q_{\theta+\delta\theta}\right)\approx(1/2)\delta\theta^{\top}S\delta\theta$ (see Appendix A for details), where $S$ is the Fisher information matrix (FIM), defined as:
$S=\expectationvalue{\nabla_{\theta}\ln q_{\theta}(\mathbf{s})\nabla_{\theta}\ln q_{\theta}(\mathbf{s})^{\top}}_{q}.$ (7)
@@
where the superscript denotes the $i$-th sample. By employing a first-order Taylor series expansion of the variational free energy, we obtain:
@@
where the factor $1/\lambda$ can be integrated into the learning rate $\alpha$. Furthermore, by substituting Eq. (10) into the constraint $(1/2)\delta\theta^{\top}S\delta\theta=\epsilon$, the adaptive learning rate $\alpha$ can be determined as:
@@
Suppose that we choose the Euclidean distance in the parameter space as the constraint: $\norm{\delta\theta}_{2}^{2}=\epsilon$, the same derivation leads to the conventional gradient descent algorithm, i.e., the FIM is equivalent to a trivial identity matrix. Intuitively, steepest descent in the distribution space is more appropriate, as a minor change in the parameter space may result in significant changes in the probability distributions before and after the update. By employing the inverse of the Fisher information matrix, $S^{-1}$, as a gradient preconditioner, natural gradient updates can effectively account for the local curvature of the manifold, leading to more efficient and stable optimization. However, the issue associated with the natural gradient method lies in the necessity of inverting the FIM, a procedure with a computational complexity of $\mathcal{O}(N_{p}^{3})$. This cubic complexity can become prohibitively expensive as the model size increases, particularly for deep neural network architectures. Next, we will introduce an efficient algorithm for resolving this issue.
@@ II.3 Efficient algorithm for implementing natural gradient descent
In a recent work, Chen and Heyl introduced a novel method called minSR [20], which enables the training of neural-network quantum states with up to $10^{6}$ parameters using stochastic reconfiguration, a quantum generalization of natural gradient method. Then Rende et al. reformulated this approach using a linear algebra identity [21]. Specifically, in the context of VAN, Eq. (4) can be estimated using $N_{b}$ samples drawn from the variational distribution:
$\eta=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}R(\mathbf{s}^{i})\nabla_{\theta}\ln q_{\theta}(\mathbf{s}^{i}).$ (12)
By defining $O_{ik}=(1/\sqrt{N_{b}})\partial\ln q_{\theta}(\mathbf{s}^{i})/\partial\theta_{k}$ and $R_{i}=(1/\sqrt{N_{b}})R(\mathbf{s}^{i})$, where $O$ is a $N_{b}\times N_{p}$ matrix and $R$ is a $N_{b}$-dimensional vector, Eq. (12) can be rewritten in matrix notation as:
@@
By leveraging a linear algebra identity: $(O^{\top}O)^{-1}O^{\top}=O^{\top}(OO^{\top})^{-1}$, we have
$\delta\theta=-\alpha O^{\top}(OO^{\top})^{-1}R,$ (16)
which exhibits a computational complexity of only $\mathcal{O}(N_{b}^{3}+N_{p}N_{b}^{2})$ compared to $\mathcal{O}(N_{p}^{3})$.
It is noteworthy that Eq. (16) is also known as natural gradient with the Moore-Penrose pseudo-inverse and has been extensively studied in the machine learning literature [29, 30, 31, 32, 33]. In the implementation, $OO^{\top}$ can be ill-conditioned and it requires regularization scheme. Usually a non-negative damping factor $\xi$ is introduced $(OO^{\top}+\xi I)^{-1}$ for numerical stability. See Ref. [34] for more details. Furthermore, this technique relies on the ability to represent the gradient as a linear combination, specifically in the form $\eta=O^{\top}R$. If the gradients do not conform to this particular structure, an alternative methodology based on Cholesky decomposition, as outlined in Ref. [22], can be employed.
@@ III Results
Figure 1: For the Sherrington-Kirkpatrick (SK) model with $N=30$ spins at the spin-glass phase transition point $\beta=1.0$, we compare the evolution of the variational free energy as training epochs and training time obtained by Adam optimizer (conventional VAN) and natural gradient (ng-VAN), using Neural Autoregressive Distribution Estimator (NADE), Masked Autoencoder for Distribution Estimation (MADE), and Transformer model as autoregressive models. The results are averaged over 10 independent parameter initialization on the same SK instance.
@@
Figure 2: For the SK model with $N=30$ spins, the relative error between the exact and variational results as a function of inverse temperature $\beta$ is presented for Adam optimizer (conventional VAN) and natural gradient (ng-VAN), using NADE, MADE, and Transformer. The inset plots show the variance of the variational free energy, which also measures the training quality. The results are averaged over 10 independent SK instances with different couplings.
@@
In our experiments, we utilized the Neural Autoregressive Distribution Estimator (NADE) [35], the Masked Autoencoder for Distribution Estimation (MADE) [36] and a decoder-only transformer [9] as our variational neural network ansätze. A concise overview of these models is provided in Appendix B. We implemented the natural gradient method using PyTorch [37] and tested it on an NVIDIA GPU. The hyperparameters of our experiments are summarized in Appendix C. We performed a series of experiments on the well-established Sherrington-Kirkpatrick (SK) model [38] as well as on the spin glasses defined on the random regular graph (RRG). In the SK model, the couplings $J_{ij}=J_{ji}$ are drawn from a Gaussian distribution with a mean of zero and a variance of $1/\sqrt{N}$. For the RRG model, we set the degree to be $d=3$, and the coupling $J_{ij}$ are randomly sampled from $\{-1,+1\}$ with equal probability.
III.1 Performance on spin glasses models

### v2

@@ Abstract
Estimating free energy is a fundamental problem in statistical mechanics. Recently, machine-learning-based methods, particularly the variational autoregressive networks (VANs) have been proposed to minimize variational free energy and to approximate the Boltzmann distribution. VAN enjoys notable advantages, including the exact computation of the normalized joint distribution and fast sampling, which are critical features often missing in Markov chain Monte Carlo algorithms. However, VAN also faces significant computational challenges. These include difficulties in the optimization of variational free energy in a complicated parameter space and slow convergence of learning. In this work, we introduce an optimization technique based on natural gradients to the VAN framework, namely ng-VAN, to enhance the learning efficiency and accuracy of the conventional VAN. The method has computational complexity cubic in the batch size rather than in the number of model parameters, hence it can be efficiently implemented for a large VAN model. We carried out extensive numerical experiments on the Sherrington-Kirkpatrick model, spin glasses on random graphs, and the two-dimensional Ising model. Our results indicate that compared with the conventional VAN, ng-VAN significantly improves the accuracy in estimating free energy and converges much faster with shorter learning time. This allows extending the VAN framework’s applicability to challenging statistical mechanics problems that were previously not accessible.
@@ I Introduction
Recently, a prominent application of machine learning methods, variational autoregressive networks (VANs) [1, 2, 3, 4], have been introduced to solve statistical mechanics problems. In VAN, an autoregressive neural network is employed as a variational ansatz $q_{\theta}(\mathbf{s})$ to approximate the Boltzmann distribution ${p_{\mathrm{eq}}(\mathbf{s})}$. Autoregressive neural networks belong to the broader category of generative models in machine learning [5] and have emerged as a powerful framework for modeling high-dimensional complex data distributions [6]. They utilize the chain rule of probabilities to parameterize the joint distribution of sequential data, such as text and images. Notable examples include large language models trained on large corpora [7, 8] and image generation by PixelCNN [9, 10]. In contrast, the learning process of VAN is no longer based on maximum likelihood but is akin to reinforcement learning, with an objective of minimizing the Kullback-Leibler divergence $D_{\mathrm{KL}}\left(q_{\theta}\middle\|p_{\mathrm{eq}}\right)$, which necessitates sampling from the variational distribution itself. The autoregressive characteristics of the model enable tractable likelihood and fast sampling [11, 12, 9, 10], a significant advantage over traditional Markov chain Monte Carlo (MCMC) methods. Moreover, VAN can provide an asymptotically unbiased estimator of physical observables [2]. Likewise, autoregressive models have also been employed to represent variational wave functions in studies of variational Monte Carlo (VMC) [13, 14, 15].
@@
Despite its advantages, VAN also faces considerable challenges, particularly regarding optimization difficulties associated with nonconvex parameter landscapes. VAN typically relies on first-order gradient-based optimization algorithms. Although state-of-the-art optimizers such as Adam [16] incorporate momentum information and show superior performances compared to traditional stochastic gradient descent [17], they may still encounter difficulties in high-dimensional parameter space and exhibit slow convergence speed. This limitation necessitates the need for more advanced optimization methods. One promising approach for solving this issue is the natural gradient descent [18, 19, 20], which has already proven effective in reinforcement learning [21, 22] as well as in VMC scenarios [23, 24]. The fundamental concept behind the natural gradient is to account for the curvature of the loss function by employing a preconditioner to rescale the gradient updates. Geometrically, the natural gradient represents the steepest descent direction in the model’s distribution space rather than in the Euclidean parameter space. However, this method requires the inversion of the Fisher information matrix, a computational complexity of $\mathcal{O}({N_{p}^{3}})$ with $N_{p}$ variational parameters. In recent advancements in VMC [25, 26] and machine learning [27], researchers have proposed efficient algorithms for implementing the natural gradient, especially in situations where the number of samples or batch size $N_{b}$ is significantly smaller than $N_{p}$. Rather than inverting the original $N_{p}\times N_{p}$ matrix, it only requires the inversion of a significantly smaller $N_{b}\times N_{b}$ matrix which shares the same spectrum properties as the Fisher information matrix. In this way, the natural gradient method can be applied to neural network models with up to $10^{6}$ parameters, which is far beyond traditional approaches.
@@
In this work, we propose to integrate the natural gradient into the VAN framework, namely ng-VAN, to enhance both learning efficiency and accuracy compared to the conventional VAN. In particular, we employ an efficient method for computing the inverse of the Fisher information matrix with computational complexity growing cubic in the batch size, rather than the number of parameters. Our results demonstrate that the natural gradient method can significantly accelerate the convergence of VAN during training and achieve more accurate estimated variational free energy compared to first-order optimizers such as Adam. The rest of the paper is organized as follows: in Sec. II, we first introduce VAN and its learning, followed by an introduction to the natural gradient method and how to implement it efficiently in the context of VAN learning. In Sec. III, we present the experimental results, and finally, in Sec. IV, we summarize our conclusions and discuss the potential applications of our methods.
@@ II.1 Variational autoregressive networks
The VANs proposed in Ref. [1] employ a variational ansatz for the target Boltzmann distribution, parameterized by a set of learnable parameters $\theta=\{\theta_{i}\}_{i=1}^{N_{p}}$, where $N_{p}$ denotes the total number of parameters. This variational distribution, denoted as $q_{\theta}(\mathbf{s})$, obeys the autoregressive property, i.e., can be decomposed into the product of conditional probabilities:
@@
$F_{q}(\theta)=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}\left[E(\mathbf{s})+\frac{1}{\beta}\ln q_{\theta}(\mathbf{s})\right].$ (3)
Moreover, the variational free energy $F_{q}(\theta)$ serves as an upper bound for the exact free energy. A crucial component required for gradient-based optimization is the gradient of the variational free energy with respect to $\theta$. Denoting $R(\mathbf{s})=E(\mathbf{s})+(1/\beta)\ln q_{\theta}(\mathbf{s})$, the gradient can be expressed as
$\eta=\nabla_{\theta}F_{q}(\theta)=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}\left[R(\mathbf{s})\nabla_{\theta}\ln q_{\theta}(\mathbf{s})\right].$ (4)
A baseline-corrected form can also be employed for variance reduction, given by $\eta=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}[(R(\mathbf{s})-b)\nabla_{\theta}\ln q_{\theta}(\mathbf{s})]$, where $b=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}[R(\mathbf{s})]$ is a constant termed as baseline. This gradient expression is known as the score function gradient estimator [28], also referred to as the REINFORCE algorithm [29], a specific instance of the policy gradient method from reinforcement learning. Note that the gradient is $\nabla_{\theta}\ln q_{\theta}(\mathbf{s})$ weighted by reward signal $R(\mathbf{s})$, indicating that configurations with large rewards are penalized in the optimization process, hence reduce the variational free energy.
VANs have shown improved performance over traditional variational mean-field methods and have attracted significant research interest, including their integration with importance sampling [2], MCMC methods [2, 30, 31], and their extension to larger sparse graphs [4]. However, as illustrated in Sec. I, it also encounters several significant challenges. In contrast to the case where a dataset is available for maximum likelihood estimation, in the training of VAN, gradient estimation in each iteration requires drawing samples from the current variational distribution. Due to the sequential nature of autoregressive sampling, this process is highly time consuming. Furthermore, gradients derived from samples based on a suboptimal variational distribution from the preceding step may not provide sufficient information for the subsequent optimization, leading to poor overall performance.
@@ II.2 Natural gradient descent
Here we present a concise and self-contained derivation of the natural gradient within the framework of optimizing VAN. The natural gradient can be derived by identifying the update $\delta\theta$ that minimizes the variational free energy $F_{q}(\theta+\delta\theta)$, subject to the constraint that the KL divergence between the current variational distribution $q_{\theta}$ and the updated distribution $q_{\theta+\delta\theta}$ is equal to a specified value $\epsilon$. Mathematically, this is formulated as
@@
The Lagrange function for this optimization problem is given by
@@
The KL divergence can be approximated by $D_{\mathrm{KL}}\left(q_{\theta}\middle\|q_{\theta+\delta\theta}\right)\approx(1/2)\delta\theta^{\top}S\delta\theta$ (see Appendix A for details), where $S$ is the Fisher information matrix (FIM), defined as
$S=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}\left[\nabla_{\theta}\ln q_{\theta}(\mathbf{s})\nabla_{\theta}\ln q_{\theta}(\mathbf{s})^{\top}\right].$ (7)
@@
where the superscript denotes the $i$-th sample. By employing a first-order Taylor series expansion of the variational free energy, we obtain
@@
where the factor $1/\lambda$ can be integrated into the learning rate $\alpha$. Furthermore, by substituting Eq. (10) into the constraint $(1/2)\delta\theta^{\top}S\delta\theta=\epsilon$, the adaptive learning rate $\alpha$ can be determined as
@@
Suppose that we choose the Euclidean distance in the parameter space as the constraint: $\norm{\delta\theta}_{2}^{2}=\epsilon$, the same derivation leads to the conventional gradient descent algorithm, i.e., the FIM is equivalent to a trivial identity matrix (see Appendix A for derivations). Intuitively, the steepest descent in the distribution space is more appropriate, as a minor change in the parameter space may result in significant changes in the probability distributions before and after the update. By employing the inverse of the Fisher information matrix, $S^{-1}$, as a gradient preconditioner, natural gradient updates can effectively account for the local curvature of the manifold, leading to more efficient and stable optimization. However, the issue associated with the natural gradient method lies in the necessity of inverting the FIM, a procedure with a computational complexity of $\mathcal{O}(N_{p}^{3})$. This cubic complexity can become prohibitively expensive as the model size increases, particularly for deep neural network architectures. Next, we will introduce an efficient algorithm for resolving this issue.
@@ II.3 Efficient algorithm for implementing natural gradient descent
In a recent work, Chen and Heyl introduced a novel method called minSR [25], which enables the training of neural-network quantum states [24] with up to $10^{6}$ parameters using stochastic reconfiguration, a quantum generalization of natural gradient method. Then Rende et al. reformulated this approach using a linear algebra identity [26]. Specifically, in the context of VAN, Eq. (4) can be estimated using $N_{b}$ samples drawn from the variational distribution:
$\eta$ $=\mathbb{E}_{\mathbf{s}\sim q_{\theta}}\left[R(\mathbf{s})\nabla_{\theta}\ln q_{\theta}(\mathbf{s})\right]$
$=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}R(\mathbf{s}^{i})\nabla_{\theta}\ln q_{\theta}(\mathbf{s}^{i}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{where}\leavevmode\nobreak\ \mathbf{s}^{i}\sim q_{\theta}(\mathbf{s})$ (12)
By defining $O_{ik}=(1/\sqrt{N_{b}})\partial\ln q_{\theta}(\mathbf{s}^{i})/\partial\theta_{k}$ and $R_{i}=(1/\sqrt{N_{b}})R(\mathbf{s}^{i})$, where $O$ is a $N_{b}\times N_{p}$ matrix and $R$ is a $N_{b}$-dimensional vector, Eq. (12) can be rewritten in matrix notation as
@@
By leveraging a linear algebra identity $(O^{\top}O)^{-1}O^{\top}=O^{\top}(OO^{\top})^{-1}$, we have
$\delta\theta=-\alpha O^{\top}(OO^{\top})^{-1}R.$ (16)
If we use a fixed learning rate $\alpha$, the computational complexity of Eq. (16) is $\mathcal{O}(N_{b}^{3}+N_{p}N_{b}^{2})$ determined by the matrix inversion of $OO^{\top}$ and the matrix multiplication of $O^{\top}(OO^{\top})^{-1}R$. In contrast to directly inverting the FIM ($\mathcal{O}(N_{p}^{3})$), our algorithm is faster especially when $N_{p}\gg N_{b}$. Besides, the computational complexity for Eq. (11) is $\mathcal{O}(N_{p}N_{b})$ instead of $\mathcal{O}(N_{p}^{2})$, determined by the dot product between $\delta\theta$ and $O^{\top}R$.
It is noteworthy that Eq. (16) is also known as natural gradient with the Moore-Penrose pseudo-inverse and has been extensively studied in the machine learning literature [32, 33, 34, 35, 36]. In the implementation, $OO^{\top}$ can be ill-conditioned and it requires regularization scheme. Usually a non-negative damping factor $\xi$ is introduced $(OO^{\top}+\xi I)^{-1}$ for numerical stability. See Ref. [37] for more details. Furthermore, in addition to the linear algebra trick utilized in Ref. [26], there are alternative techniques for efficiently solving the inverse of the FIM, such as the algorithm based on Cholesky decomposition proposed in Ref. [27].
@@ III Results
Figure 1: For the Sherrington-Kirkpatrick (SK) model with $N=30$ spins at the spin-glass phase transition point $\beta=1.0$, we compare the evolution of the variational free energy as training epochs and training time obtained by Adam optimizer (conventional VAN) and natural gradient (ng-VAN), using NADE, MADE, and the transformer model as autoregressive models. The results are averaged over 10 independent parameter initialization on the same SK instance.
@@
Figure 2: For the SK model with $N=30$ spins, the relative error between the exact and variational results as a function of inverse temperature $\beta$ is presented for Adam optimizer (conventional VAN) and natural gradient (ng-VAN), using NADE, MADE, and the transformer. The inset plots show the variance of the variational free energy, which also measures the training quality. The results are averaged over 10 independent SK instances with different couplings.
@@
In our experiments, we utilized the Neural Autoregressive Distribution Estimator (NADE) [11], the Masked Autoencoder for Distribution Estimation (MADE) [12], a decoder-only transformer [7] and PixelCNN [9, 10] as our variational neural network ansätze. A concise overview of these models is provided in Appendix B. We implemented the natural gradient method using PyTorch [38] and tested it on an NVIDIA GPU. The hyperparameters of our experiments are summarized in Appendix C. We performed a series of experiments on the well-established Sherrington-Kirkpatrick (SK) model [39], spin glasses defined on the random regular graph (RRG) as well as the two-dimensional Ising model. In the SK model, the couplings $J_{ij}=J_{ji}$ are drawn from a Gaussian distribution with a mean of zero and a variance of $1/\sqrt{N}$. For the RRG model, we set the degree to be $d=3$, and the coupling $J_{ij}$ are randomly sampled from $\{-1,+1\}$ with equal probability.
III.1 Performance on variational free energy